Suppose a car can accelerate from rest to 200 km/h in 10 s. What is its acceleration?


1
Expert's answer
2021-10-24T18:20:15-0400

Given:

"v_i=0\:\rm m/s"

"v_f=200\:\rm km/hr=55.6\:\rm m/s"

"t=10\:\rm s"


The acceleration of a car

"a=\frac{v_f-v_i}{t}=\frac{55.6-0}{10}=5.56\:\rm m/s^2"